# Learning Neural Contracting Dynamics: Extended Linearization and Global Guarantees

### v2

@@ Abstract
Global stability and robustness guarantees in learned dynamical systems are essential to ensure well-behavedness of the systems in the face of uncertainty. We present Extended Linearized Contracting Dynamics (ELCD), the first neural network-based dynamical system with global contractivity guarantees in arbitrary metrics. The key feature of ELCD is a parametrization of the extended linearization of the nonlinear vector field. In its most basic form, ELCD is guaranteed to be (i) globally exponentially stable, (ii) equilibrium contracting, and (iii) globally contracting with respect to some metric. To allow for contraction with respect to more general metrics in the data space, we train diffeomorphisms between the data space and a latent space and enforce contractivity in the latent space, which ensures global contractivity in the data space. We demonstrate the performance of ELCD on the $2$D, $4$D, and $8$D LASA datasets.
Machine Learning, contraction, ICML
[…]
@@ 1 Introduction
To enforce stability guarantees, a popular approach has been to ensure that the learned dynamics admit a unique equilibrium point and have a Lyapunov function, e.g. (Manek & Kolter, 2019). While popular, approaches based on Lyapunov functions typically struggle to provide general robustness guarantees since global asymptotic stability does not ensure robustness guarantees in the presence of disturbances. Indeed, input-to-state stability of global asymptotically stable dynamical systems needs to be separately established, e.g., see Chapter 5 in (Khalil, 2002).
To ensure robustness in addition to stability, there has been increased interest in learning contracting dynamics. A dynamical system is said to be contracting if any two trajectories converge to one another exponentially quickly with respect to some metric (Lohmiller & Slotine, 1998). It is known that if a system is contracting, it is also exponentially incrementally input-to-state stable (Tsukamoto et al., 2021). Additionally, if the system is autonomous, it admits a unique equilibrium, is input-to-state stable, and has two Lyapunov functions that establish exponential stability of the equilibrium (Bullo, 2023). Since establishing contractivity globally requires satisfying a matrix partial differential inequality everywhere, prior works have focused on establishing contractivity in the neighborhood of training data.
@@ 1.1 Related Works
Stable, but not necessarily contracting dynamics. Numerous works have aimed to learn stable dynamical systems from data including (Khansari-Zadeh & Billard, 2011; Manek & Kolter, 2019; Urain et al., 2020; Rana et al., 2020; Zhi et al., 2022; Bevanda et al., 2022; Fan et al., 2022). In (Khansari-Zadeh & Billard, 2011), the authors introduce the Stable Estimator of Dynamical Systems (SEDS), which leverages a Gaussian mixture model to approximate the dynamics and enforce asymptotic stability via constraints on learnable parameters. In (Manek & Kolter, 2019), the authors jointly learn the dynamics and a Lyapunov function for the system and project the dynamics onto the set of dynamics which enforce an exponentially decay the Lyapunov function. This projection is done in closed-form and establishes global exponential convergence of the dynamics to the origin. In (Urain et al., 2020), the authors introduce Imitation Flow where trajectories are mapped to a latent space where states evolve in time according to a stable SDE. In (Rana et al., 2020), Euclideanizing Flows is introduced where the latent dynamics are enforced to follow natural gradient dynamics (Amari, 1998). A similar approach is taken in (Zhi et al., 2022) where additionally collision avoidance is considered.
@@
Existing works on learning contracting dynamics. Learning contracting vector fields from demonstrations has attracted attention due to robustness guarantees (Ravichandar et al., 2017; Sindhwani et al., 2018; Sun et al., 2021; Tsukamoto et al., 2021; Beik-Mohammadi et al., 2024). In (Ravichandar et al., 2017), the dynamics are defined using a Gaussian mixture model and the contraction metric is parametrized to be a symmetric matrix of polynomial functions. Convergence to an equilibrium is established using partial contraction theory (Wang & Slotine, 2005). In (Sindhwani et al., 2018), the dynamics are defined via an optimization problem over a reproducing kernel Hilbert space; the dynamics are constrainted to be locally contracting around the data points, i.e., there are points in state space where the dynamics are not contracting. In (Sun et al., 2021) and (Tsukamoto et al., 2021), the authors study controlled dynamical systems of the form $\dot{x}=f(x,u)$, where $u$ is a control input and train neural networks to find a feedback controller such that the closed-loop dynamics are approximately contracting, i.e., contractivity is not enforced but instead lack of contractivity is penalized in the cost function.
@@
Recent work, (Beik-Mohammadi et al., 2024), has proposed a Neural Contractive Dynamical System (NCDS) which learns a dynamical system which is explicitly contracting to a trajectory. NCDS parameterizes contracting vector fields by learning symmetric Jacobians and performing line integrals to evaluate the underlying vector field. NCDS is computationally costly because of this integration. Additionally, the Jacobian parametrization used is overly restrictive, because not all contracting vector fields have symmetric Jacobians. Constraining the vector field to have symmetric Jacobian is equivalent to enforcing that the vector field is a negative gradient flow and contracting with respect to the identity metric (Wensing & Slotine, 2020). For scalability to higher dimensions, NCDS leverages a latent space structure where an encoder maps the data space to a lower-dimensional space and enforces contracting dynamics in this latent space. Then a decoder “projects” the latent-space dynamics to the full data space. It is then argued that on the submanifold defined by the image of the decoder, the dynamics are contracting.
@@ 1.2 Contributions
Our example in Section 3.4 provides theoretical justification for why the diffeomorphism and the learned contracting model must be trained jointly. In summary, if the diffeomorphism is trained first, and transforms the data to a latent space, the model class may not be expressive enough to accurately represent the true latent dynamics. If the diffeomorphism and dynamics are trained simultaneously, this limitation is overcome. This is in contrast to (Beik-Mohammadi et al., 2024) which trains the diffeomorphism independently as a variational autoencoder and then trains the model after on the transformed data.
Our model, ELCD, directly improves on NCDS in several ways. We parameterize the vector field directly instead of parameterizing its Jacobian. Doing so prevents us from needing to integrate the Jacobian to calculate the vector field and thus speeds up training and inference. ELCD is also more expressive than NCDS because it can represent vector fields with asymmetric Jacobians. Additionally, ELCD is guaranteed to be contracting to an arbitrary equilibrium point, either selected or learned, at all training steps. NCDS, in contrast, must learn the equilibrium point over the course of training. Additionally, NCDS learns an encoder and decoder for a lower-dimensional latent space and thus can only be contracting on the submanifold defined by the image of the decoder. From initial conditions that are not on this submanifold, NCDS may not exhibit contracting behavior. In contrast, since the latent space of ELCD is of the same dimension as the data space, we train diffeomorphisms that ensure global contractivity in the data space. ELCD exhibits better performance than NCDS at reproducing trajectories from the LASA (Lemme et al., 2015) dataset, including achieving 45% less normalized time warping in 37% less training time for 8D-LASA.
@@ 2 Preliminaries
A contracting dynamical system is one for which any two trajectories converge exponentially quickly. From (Lohmiller & Slotine, 1998), for a continuously differentiable map $f:^{d}\rightarrow^{d}$, the dynamical system $\dot{x}=f(x)$ is contracting with rate $c>0$ if there exists a continuously-differentiable matrix-valued map $M:^{d}\rightarrow^{d\times d}$ and two constants $a_{0},a_{1}>0$ such that for all $x\in^{n}$, $M(x)=M(x)^{\top}$ and $a_{0}I_{d}\succeq M(x)\succeq a_{1}I_{d}$ and additionally satisfies for all $x$
@@
The map $M$ is called the contraction metric and the notation $\dot{M}(x)$ is shorthand for the $n\times n$ matrix whose $(i,j)$ entry is $\dot{M}(x)_{ij}=\nabla M_{ij}(x)^{\top}f(x)$. A central result in contraction theory is that any dynamical system $\dot{x}=f(x)$ satisfying (2) has any two trajectories converging to one another exponentially quickly (Lohmiller & Slotine, 1998; Manchester & Slotine, 2017; Tsukamoto et al., 2021). Specifically, there exists $L\geq 1$ such that any two trajectories $x_{1},x_{2}$ of the dynamical system satisfy
@@
In other words, contractivity establishes exponential incremental stability (Angeli, 2002).
We note that any contracting dynamical system enjoys numerous useful properties including the existence of a unique exponentially stable equilibrium and exponential input-to-state stability when perturbed by a disturbance. We refer to (Lohmiller & Slotine, 1998) for more useful properties of contracting dynamical systems and (Bullo, 2023) for a recent monograph on the subject.
@@ 2.1 Contracting Linear Systems
which implies that $A$ has all eigenvalues with $\mathrm{Re}(\lambda(A))\leq-c$, see Theorem 8.2 in (Hespanha, 2009). In other words, for linear systems, contractivity is equivalent to stability.
Although the condition (4) is convex in $M$ at fixed $A$, the task of learning both $(A,M)$ simultaneously from data is not (jointly) convex. Instead, different methods must be employed such as alternating minimization. A similar argument is made in (Manek & Kolter, 2019) in the context of stability and here we show that the same is true of contractivity.
@@ 2.2 Contractivity and Exponential Stability for Nonlinear Systems
One advantage of learning asymptotically stable dynamics compared to contracting ones is that the function $V$ is scalar-valued, while the contraction metric $M$ is matrix-valued. A disadvantage of learning asymptotically stable dynamics compared to contracting ones is that we are required to know the location of the equilibrium point beforehand and no robustness property of the learned dynamics is automatically enforced. To this end, there are works in the literature that enforce an equilibrium point at the origin and learn the dynamics and/or the Lyapunov function under this assumption (Manek & Kolter, 2019). In the case of contractivity, existence and uniqueness of an equilibrium point is implied by the condition (2) and it can be directly parametrized to best suit the data.
@@ 3.1 Motivation
Lemma 2 (Mean-value theorem (Proposition 2.4.7 in (Abraham et al., 1988))).
@@
This factorization is referred to as extended linearization in analogy to standard linearization, i.e., for $x$ close to $x^{*}$, $f(x)\approx D\mkern-0.75mu{f}(x^{*})(x-x^{*})$. We remark that when $d\geq 2$, extended linearization is not unique, and for a given $f$, there may exist several $A$ such that $f(x)=A(x,x^{*})(x-x^{*})$. In other words, (8) showcases one valid choice for $A$ such that this factorization holds. Indeed, this nonuniqueness has been leveraged in some prior words, e.g. Section 3.1.3 in (Tsukamoto et al., 2021), to yield less conservative contractivity conditions.
Extended linearization is also sometimes referred to as apparent linearization and the state-dependent coefficient parametrization and has a rich history in nonlinear control theory, specifically in state-dependent linear-quadratic regulator (LQR) design; see the survey (Çimen, 2008). Moreover, it was recently demonstrated in (Richards et al., 2023) that enforcing this extended linearization structure during learning can be more data-efficient.
@@ 3.2 Parametrization of $A(x,x^{*})$
We prove that a vector field parametrized this way is guaranteed to be (i) globally exponentially stable, (ii) equilibrium contracting as defined in (Davydov et al., 2022), and (iii) contracting in some metric. The key tools are partial contraction theory (Wang & Slotine, 2005) and a converse contraction theorem.
@@
We use the method of partial contraction as in (Wang & Slotine, 2005). To this end, define the virtual system
@@
Clearly, the bound (12) implies exponential convergence of trajectories of the dynamical system (10) to $x^{*}$. Moreover, this bound exactly characterizes equilibrium contractivity, as was defined in (Davydov et al., 2022). Notably, using the language of logarithmic norms, Theorem 33 in (Davydov et al., 2022) establishes a similar result to Theorem 3 without invoking a virtual system.
@@
admits a unique solution for each $x$ (Giesl et al., 2021). In other words, the unique $M$ solving this matrix PDE serves as the contraction metric and will satisfy (2) with suitable choice for $c$. The following proposition provides the explicit solution for $M$ in terms of the solution to the matrix PDE.
Proposition 4 (Theorem 2.8 in (Giesl et al., 2021)).
@@
While it is challenging, in general, to compute the metric (15), numerical considerations for approximating it arbitrarily closely are presented in (Giesl et al., 2023). In practice, one would select $C(x)=I_{d}$ and evaluate all integrals numerically.
We remark that our parametrization for $A$ in (11) is similar to the parametrization for the Jacobian of $f$ that was presented in (Beik-Mohammadi et al., 2024). There are a few key differences. Notably, $A(x,x^{*})$ may be asymmetric while the Jacobian in equation (3) in (Beik-Mohammadi et al., 2024) is always symmetric. Notably, since their Jacobian is symmetric and negative definite, the vector field $\dot{x}=f(x)$ is a negative gradient flow, $\dot{x}=-\nabla V(x)$, for some strongly convex $V$. On the contrary, our dynamics (10) can exhibit richer behaviors than negative gradient flows in view of the asymmetry in $A(x,x^{*})$. Additionally, since the dynamics in (Beik-Mohammadi et al., 2024) with their parametrization can be represented as $\dot{x}=-\nabla V(x)$ for some strongly convex $V$, it is guaranteed to be contracting in the identity metric, $M(x)=I_{d}$ (Wensing & Slotine, 2020). On the other hand, our dynamics (10) is not necessarily contracting in the identity metric and instead is contracting in a more complex metric given in (15).
@@ 3.3 Latent Space Representation
Latent space learning is possible because contraction is invariant under differential coordinate changes. From Theorem 2 of (Manchester & Slotine, 2017), given a dynamical system $\dot{x}=f(x)$, $f:^{d}\to^{d}$, with $f$ satisfying (1), the system will also be contracting under the coordinate change $z=\phi(x)$ if $\phi:^{d}\rightarrow^{d}$ is a smooth diffeomorphism. Specifically, if $\dot{x}=f(x)$ is contracting with metric $M$, then the system that evolves $z$ is contracting as well with metric given by $\tilde{M}(x)=D\mkern-0.75mu{\phi}(x)^{-\top}M(x)D\mkern-0.75mu{\phi}(x)^{-1}$, where $z=\phi(x)$ and $D\mkern-0.75mu{\phi}$ is the Jacobian of the coordinate transform. In other words, We can learn vector fields that are contracting in an arbitrary metric by training a vector field $f$ which is parametrized as (10) and using a coordinate transform $\phi$.
(Beik-Mohammadi et al., 2024) treats the coordinate transform as a Variational Autoencoder (VAE) (Kingma & Welling, 2014). Their training procedure consists of two steps: first training the coordinate transform with VAE training, then training the function $f$ in the new learned coordinates.

### v3

@@ Abstract
Global stability and robustness guarantees in learned dynamical systems are essential to ensure well-behavedness of the systems in the face of uncertainty. We present Extended Linearized Contracting Dynamics (ELCD), the first neural network-based dynamical system with global contractivity guarantees in arbitrary metrics. The key feature of ELCD is a parametrization of the extended linearization of the nonlinear vector field. In its most basic form, ELCD is guaranteed to be (i) globally exponentially stable, (ii) equilibrium contracting, and (iii) globally contracting with respect to some metric. To allow for contraction with respect to more general metrics in the data space, we train diffeomorphisms between the data space and a latent space and enforce contractivity in the latent space, which ensures global contractivity in the data space. We demonstrate the performance of ELCD on the high dimensional LASA, multi-link pendulum, and Rosenbrock datasets.
@@ 1 Introduction
To enforce stability guarantees, a popular approach has been to ensure that the learned dynamics admit a unique equilibrium point and have a Lyapunov function, e.g. [27]. While popular, approaches based on Lyapunov functions typically struggle to provide general robustness guarantees since global asymptotic stability does not ensure robustness guarantees in the presence of disturbances. Indeed, input-to-state stability of global asymptotically stable dynamical systems needs to be separately established, e.g., see Chapter 5 in [20].
To ensure robustness in addition to stability, there has been increased interest in learning contracting dynamics. A dynamical system is said to be contracting if any two trajectories converge to one another exponentially quickly with respect to some metric [25]. It is known that if a system is contracting, it is also exponentially incrementally input-to-state stable [34]. Additionally, if the system is autonomous, it admits a unique equilibrium, is input-to-state stable, and has two Lyapunov functions that establish exponential stability of the equilibrium [9]. Since establishing contractivity globally requires satisfying a matrix partial differential inequality everywhere, prior works have focused on establishing contractivity in the neighborhood of training data.
@@ 1.1 Related Works
Stable, but not necessarily contracting dynamics. Numerous works have aimed to learn stable dynamical systems from data including [21, 27, 35, 29, 38, 7, 14]. In [21], the authors introduce the Stable Estimator of Dynamical Systems (SEDS), which leverages a Gaussian mixture model to approximate the dynamics and enforce asymptotic stability via constraints on learnable parameters. In [27], the authors jointly learn the dynamics and a Lyapunov function for the system and project the dynamics onto the set of dynamics which enforce an exponentially decay the Lyapunov function. This projection is done in closed-form and establishes global exponential convergence of the dynamics to the origin. In [35], the authors introduce Imitation Flow where trajectories are mapped to a latent space where states evolve in time according to a stable SDE. In [29], Euclideanizing Flows is introduced where the latent dynamics are enforced to follow natural gradient dynamics [2]. A similar approach is taken in [38] where additionally collision avoidance is considered.
@@
Existing works on learning contracting dynamics. Learning contracting vector fields from demonstrations has attracted attention due to robustness guarantees [30, 32, 33, 34, 5]. In [30], the dynamics are defined using a Gaussian mixture model and the contraction metric is parametrized to be a symmetric matrix of polynomial functions. Convergence to an equilibrium is established using partial contraction theory [36]. In [32], the dynamics are defined via an optimization problem over a reproducing kernel Hilbert space; the dynamics are constrainted to be locally contracting around the data points, i.e., there are points in state space where the dynamics are not contracting. In [33] and [34], the authors study controlled dynamical systems of the form $\dot{x}=f(x,u)$, where $u$ is a control input and train neural networks to find a feedback controller such that the closed-loop dynamics are approximately contracting, i.e., contractivity is not enforced but instead lack of contractivity is penalized in the cost function.
@@
Recent work, [5], has proposed a Neural Contractive Dynamical System (NCDS) which learns a dynamical system which is explicitly contracting to a trajectory. NCDS parameterizes contracting vector fields by learning symmetric Jacobians and performing line integrals to evaluate the underlying vector field. NCDS is computationally costly because of this integration. Additionally, the Jacobian parametrization used is overly restrictive, because not all contracting vector fields have symmetric Jacobians. Constraining the vector field to have symmetric Jacobian is equivalent to enforcing that the vector field is a negative gradient flow and contracting with respect to the identity metric [37]. For scalability to higher dimensions, NCDS leverages a latent space structure where an encoder maps the data space to a lower-dimensional space and enforces contracting dynamics in this latent space. Then a decoder “projects" the latent-space dynamics to the full data space. It is then argued that on the submanifold defined by the image of the decoder, the dynamics are contracting.
@@ 1.2 Contributions
Our example in Section 3.4 provides theoretical justification for why the diffeomorphism and the learned contracting model must be trained jointly. In summary, if the diffeomorphism is trained first, and transforms the data to a latent space, the model class may not be expressive enough to accurately represent the true latent dynamics. If the diffeomorphism and dynamics are trained simultaneously, this limitation is overcome. This is in contrast to [5] which trains the diffeomorphism independently as a variational autoencoder and then trains the model after on the transformed data.
Our model, ELCD, directly improves on NCDS in several ways. We parameterize the vector field directly instead of parameterizing its Jacobian. Doing so prevents us from needing to integrate the Jacobian to calculate the vector field and thus speeds up training and inference. ELCD is also more expressive than NCDS because it can represent vector fields with asymmetric Jacobians. Additionally, ELCD is guaranteed to be contracting to an arbitrary equilibrium point, either selected or learned, at all training steps. NCDS, in contrast, must learn the equilibrium point over the course of training. Additionally, NCDS learns an encoder and decoder for a lower-dimensional latent space and thus can only be contracting on the submanifold defined by the image of the decoder. From initial conditions that are not on this submanifold, NCDS may not exhibit contracting behavior. In contrast, since the latent space of ELCD is of the same dimension as the data space, we train diffeomorphisms that ensure global contractivity in the data space. ELCD exhibits better performance than NCDS at reproducing trajectories from the LASA [24], n-link Pendulum, and Rosenbrock datasets. We additionally compare against the Euclideanizing Flow [29], and Stable Deep Dynamics [27] models.
@@ 2 Preliminaries
A contracting dynamical system is one for which any two trajectories converge exponentially quickly. From [25], for a continuously differentiable map $f:{}^{d}\rightarrow{}^{d}$, the dynamical system $\dot{x}=f(x)$ is contracting with rate $c>0$ if there exists a continuously-differentiable matrix-valued map $M:{}^{d}\rightarrow{}^{d\times d}$ and two constants $a_{0},a_{1}>0$ such that for all $x\in{}^{n}$, $M(x)=M(x)^{\top}$ and $a_{0}I_{d}\succeq M(x)\succeq a_{1}I_{d}$ and additionally satisfies for all $x$
@@
The map $M$ is called the contraction metric and the notation $\dot{M}(x)$ is shorthand for the $n\times n$ matrix whose $(i,j)$ entry is $\dot{M}(x)_{ij}=\nabla M_{ij}(x)^{\top}f(x)$. A central result in contraction theory is that any dynamical system $\dot{x}=f(x)$ satisfying (2) has any two trajectories converging to one another exponentially quickly [25, 26, 34]. Specifically, there exists $L\geq 1$ such that any two trajectories $x_{1},x_{2}$ of the dynamical system satisfy
@@
In other words, contractivity establishes exponential incremental stability [4].
We note that any contracting dynamical system enjoys numerous useful properties including the existence of a unique exponentially stable equilibrium and exponential input-to-state stability when perturbed by a disturbance. We refer to [25] for more useful properties of contracting dynamical systems and [9] for a recent monograph on the subject.
@@ 2.1 Contracting Linear Systems
which implies that $A$ has all eigenvalues with $\mathrm{Re}(\lambda(A))\leq-c$, see Theorem 8.2 in [17]. In other words, for linear systems, contractivity is equivalent to stability.
Although the condition (4) is convex in $M$ at fixed $A$, the task of learning both $(A,M)$ simultaneously from data is not (jointly) convex. Instead, different methods must be employed such as alternating minimization. A similar argument is made in [27] in the context of stability and here we show that the same is true of contractivity.
@@ 2.2 Contractivity and Exponential Stability for Nonlinear Systems
One advantage of learning asymptotically stable dynamics compared to contracting ones is that the function $V$ is scalar-valued, while the contraction metric $M$ is matrix-valued. A disadvantage of learning asymptotically stable dynamics compared to contracting ones is that we are required to know the location of the equilibrium point beforehand and no robustness property of the learned dynamics is automatically enforced. To this end, there are works in the literature that enforce an equilibrium point at the origin and learn the dynamics and/or the Lyapunov function under this assumption [27]. In the case of contractivity, existence and uniqueness of an equilibrium point is implied by the condition (2) and it can be directly parametrized to best suit the data.
@@ 3.1 Motivation
Lemma 2 (Mean-value theorem (Proposition 2.4.7 in [1])).
@@
This factorization is referred to as extended linearization in analogy to standard linearization, i.e., for $x$ close to $x^{*}$, $f(x)\approx D\mkern-0.75mu{f}(x^{*})(x-x^{*})$. We remark that when $d\geq 2$, extended linearization is not unique, and for a given $f$, there may exist several $A$ such that $f(x)=A(x,x^{*})(x-x^{*})$. In other words, (8) showcases one valid choice for $A$ such that this factorization holds. Indeed, this nonuniqueness has been leveraged in some prior words, e.g. Section 3.1.3 in [34], to yield less conservative contractivity conditions.
@@ 3.2 Parametrization of $A(x,x^{*})$
We prove that a vector field parametrized this way is guaranteed to be (i) globally exponentially stable, (ii) equilibrium contracting as defined in [10], and (iii) contracting in some metric. The key tools are partial contraction theory [36] and a converse contraction theorem.
@@
We use the method of partial contraction as in [36]. To this end, define the virtual system
@@
Clearly, the bound (12) implies exponential convergence of trajectories of the dynamical system (10) to $x^{*}$. Moreover, this bound exactly characterizes equilibrium contractivity, as was defined in [10]. Notably, using the language of logarithmic norms, Theorem 33 in [10] establishes a similar result to Theorem 3 without invoking a virtual system.
@@
admits a unique solution for each $x$ [15]. In other words, the unique $M$ solving this matrix PDE serves as the contraction metric and will satisfy (2) with suitable choice for $c$. The following proposition provides the explicit solution for $M$ in terms of the solution to the matrix PDE.
Proposition 4 (Theorem 2.8 in [15]).
@@
While it is challenging, in general, to compute the metric (15), numerical considerations for approximating it arbitrarily closely are presented in [16]. In practice, one would select $C(x)=I_{d}$ and evaluate all integrals numerically.
We remark that our parametrization for $A$ in (11) is similar to the parametrization for the Jacobian of $f$ that was presented in [5]. There are a few key differences. Notably, $A(x,x^{*})$ may be asymmetric while the Jacobian in equation (3) in [5] is always symmetric. Notably, since their Jacobian is symmetric and negative definite, the vector field $\dot{x}=f(x)$ is a negative gradient flow, $\dot{x}=-\nabla V(x)$, for some strongly convex $V$. On the contrary, our dynamics (10) can exhibit richer behaviors than negative gradient flows in view of the asymmetry in $A(x,x^{*})$. Additionally, since the dynamics in [5] with their parametrization can be represented as $\dot{x}=-\nabla V(x)$ for some strongly convex $V$, it is guaranteed to be contracting in the identity metric, $M(x)=I_{d}$ [37]. On the other hand, our dynamics (10) is not necessarily contracting in the identity metric and instead is contracting in a more complex metric given in (15).
@@ 3.3 Latent Space Representation
Latent space learning is possible because contraction is invariant under differential coordinate changes. From Theorem 2 of [26], given a dynamical system $\dot{x}=f(x)$, $f:{}^{d}\to{}^{d}$, with $f$ satisfying (1), the system will also be contracting under the coordinate change $z=\phi(x)$ if $\phi:{}^{d}\rightarrow{}^{d}$ is a smooth diffeomorphism. Specifically, if $\dot{x}=f(x)$ is contracting with metric $M$, then the system that evolves $z$ is contracting as well with metric given by $\tilde{M}(x)=D\mkern-0.75mu{\phi}(x)^{-\top}M(x)D\mkern-0.75mu{\phi}(x)^{-1}$, where $z=\phi(x)$ and $D\mkern-0.75mu{\phi}$ is the Jacobian of the coordinate transform. In other words, We can learn vector fields that are contracting in an arbitrary metric by training a vector field $f$ which is parametrized as (10) and using a coordinate transform $\phi$.
NCDS, [5], treats the coordinate transform as a Variational Autoencoder (VAE) [22]. Their training procedure consists of two steps: first training the coordinate transform with VAE training, then training the function $f$ in the new learned coordinates.
3.4 On the Interdependence of Diffeomorphism and Learned Dynamics
To demonstrate the need for a coordinate transform, we consider the task of learning a vector field that fits trajectories generated by the linear system $\dot{x}=Ax$ with $A=\begin{pmatrix}-1&4\\ 0&-1\end{pmatrix}.$ Clearly, this linear system cannot be represented in the form (10) with parametrization (11). To see this fact, we observe that $A+A^{\top}$ is not negative definite and thus no choice of $P_{s}$ or $P_{a}$ can represent this linear system. To remedy this issue, one can take the linear coordinate transform $z=\phi(x)=Px$, where $P=\begin{pmatrix}1&0\\ 0&4\end{pmatrix}$. Then the $z$-dynamics read
$\dot{z}=PAP^{-1}z=\begin{pmatrix}-1&1\\ 0&-1\end{pmatrix}z$ (17)
and now the symmetric part of $PAP^{-1}$ is negative definite and thus we can find suitable choices of $P_{s},P_{a}$. Specifically, we can take $\alpha\in(0,1/2)$, let $P_{a}(z)=\begin{pmatrix}0&0\\ 1/2&0\end{pmatrix}z$, and let $P_{s}(z)=Qz$ where $Q$ is the matrix square root of $\begin{pmatrix}1&-1/2\\ -1/2&1\end{pmatrix}-\alpha I_{2}$. Note that in this toy example, asymmetry is essential to exactly represent these dynamics in the latent space, $z$. If we used the parametrization in [5], we would not be able to represent these dynamics since they have an asymmetric Jacobian.